OpenStudy (sophiesky):

Please help, I need someone to talk me through the steps... Determine the domain of the function: 3x÷x(x^2−25)

jimthompson5910 (jim_thompson5910):

Sounds like you're dealing with \[\Large \frac{3x}{x(x^2-25)}\] Is my assumption correct?

OpenStudy (sophiesky):

Yes, your assumption is correct. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Focus strictly on the denominator: \(\Large x(x^2-25)\)

jimthompson5910 (jim_thompson5910):

Set it equal to zero and solve for x. Tell me what you get

OpenStudy (sophiesky):

Alright I'm stuck on \[x^3=25\]

jimthompson5910 (jim_thompson5910):

Did you use the idea that if A*B = 0, then A = 0 or B = 0 ?

jimthompson5910 (jim_thompson5910):

x(x^2-25) = 0 x = 0 or x^2-25 = 0 so we know that x = 0. That's one of the there values. There are 2 more Solve x^2 - 25 = 0 for x

OpenStudy (sophiesky):

Ohhhh alright one sec

OpenStudy (sophiesky):

So \[x=\sqrt{25}\]

jimthompson5910 (jim_thompson5910):

what is the square root of 25 equal to?

OpenStudy (sophiesky):

Oh 5 duh.. Sorry, it's been a long day in pre-cal :P

jimthompson5910 (jim_thompson5910):

don't forget about the plus minus eg: x^2 = 100 x = 10 or x = -10

OpenStudy (sophiesky):

Ugh.. Okay so x=0, x=5, x=-5?

jimthompson5910 (jim_thompson5910):

So x = 0 or x = 5 or x = -5 will make the denominator equal to zero. Therefore, we must kick these values out of the domain. Any other real number will be a valid input.

OpenStudy (sophiesky):

Ohhh I see, so the domain is all real numbers except for those three!

jimthompson5910 (jim_thompson5910):

Correct. In `set-builder notation` the domain would be \[\Large \left\{x| x\in\mathbb{R}, \ x \ne 0, \ x \ne 5, \ x \ne -5\right\}\] In `interval notation`, the domain would be \[\Large (-\infty, -5) \cup (-5, 0) \cup (0, 5)\cup(5,\infty)\] The idea here is that we start with (-infty, infty) and we poke holes at the values where we're pulling out of the domain

OpenStudy (sophiesky):

Got it! I think I understand now! :)

jimthompson5910 (jim_thompson5910):

I'm glad it makes more sense